please help me with this problem.
Solve the equation.
9y^2 + 12y + 4 = 0
thanks so much.

Answer by minerva(22)

9y^2 + 12y + 4 = 0
9y^2 +6y+6y+4=0
3y(3y+2) + 2(3y+2)=0
3y+2=0
y=-2/3